# Some topological properties of the intrinsic volume metric

### v1

@@ Abstract.
We prove that for fixed $j\in\{2,\ldots,d-1\}$, the $j$th intrinsic volume metric $\delta_{j}$ is not uniformly bounded, and the metric space $(\mathcal{K}_{j}^{d},\delta_{j})$ is incomplete. In particular, this proves that for $j\in\{2,\ldots,d-1\}$, the Hausdorff and $j$th intrinsic volume metrics do not induce the same topology on $\mathcal{K}_{j}^{d}$.
@@ 1. Introduction and main results
where $B_{d}$ is the $d$-dimensional Euclidean unit ball centered at the origin. The symmetric difference metric is defined by $d_{S}(K,L)=\operatorname{vol}_{d}(K\triangle L)$, where $\operatorname{vol}_{d}$ denotes the $d$-dimensional volume functional and $K\triangle L=(K\setminus L)\cup(L\setminus K)$. It is well-known that the metric spaces $(\mathcal{K}^{d},d_{H})$ and $(\mathcal{K}_{d}^{d}\cup\{\varnothing\},d_{S})$ are complete. As explained in [BH-2024], in the latter case the empty set is appended to include limits of sequences of convex bodies $\{K_{i}\}\subset\mathcal{K}^{d}_{d}$ which converge in the Hausdorff metric to a set $K_{0}\in\mathcal{K}^{d}$ with $\dim(K_{0})<d$. More specifically, if $K_{i}\to K_{0}\in\mathcal{K}^{d}$ and $\dim(K_{0})<d$, then for such a sequence we have $d_{S}(K_{i},\varnothing)=\operatorname{vol}_{d}(K_{i})\to 0$, so $K_{i}\to\varnothing$ with respect to $d_{S}$. It is also known that on uniformly bounded subsets of $\mathcal{K}_{d}^{d}$, the metrics $d_{H}$ and $d_{S}$ are equivalent. For more background on the Hausdorff and symmetric difference metrics, see the article [ShephardWebster-1965] by Shephard and Webster.
For $j\in\{1,\ldots,d\}$ and $K,L\in\mathcal{K}_{j}^{d}$, the $j$th intrinsic volume metric $\delta_{j}(K,L)$ was defined in [BH-2024] by
(1) $\delta_{j}(K,L):=\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\int_{\operatorname{Gr}(d,j)}\operatorname{vol}_{j}\left((P_{H}K)\triangle(P_{H}L)\right)d\nu_{j}(H).$
@@
Here and throughout the paper, $\operatorname{Gr}(d,j)$ is the Grassmannian manifold of all $j$-dimensional subspaces of $\mathbb{R}^{d}$, $\nu_{j}$ is the unique Haar probability measure on $\operatorname{Gr}(d,j)$, $P_{H}$ denotes the orthogonal projection of $\mathbb{R}^{d}$ into the subspace $H\in\operatorname{Gr}(d,j)$, and $\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]=\binom{d}{j}\frac{\operatorname{vol}_{d}(B_{d})}{\operatorname{vol}_{j}(B_{j})\operatorname{vol}_{d-j}(B_{d-j})}$ is the flag coefficient of Klain and Rota [Klain-Rota]. The intrinsic volume metric has found natural applications to the approximation of convex bodies by polytopes; see [BH-2024, BHK, HSW-2025].
It was conjectured in [BH-2024] that $d_{H}$ and $\delta_{j}$ induce the same topology on $\mathcal{K}_{j}^{d}$. In Open Question 2.2 of [BH-2024], they asked the following
Open Question 1.1 (uniform bound).
As remarked in [BH-2024], a positive answer to this question would imply that a sequence $\{K_{i}\}$ which converges with respect to $\delta_{j}$ also converges with respect to $d_{H}$. Moreover, as noted in [BH-2024], for the special case $j=d$, this question was answered affirmatively by Shephard and Webster in [ShephardWebster-1965, Lemma 11], and for $j=1$ see [Vitale-1985, Theorem 3] and [Florian-1989, Theorem 2].
We show that, interestingly, when $2\leq j\leq d-1$, the answers to these questions are negative.
Theorem 1.2.
@@
Theorem 1.4.
It is thus natural to ask if appending the empty set to $\mathcal{K}_{j}^{d}$ will make $(\mathcal{K}_{j}^{d}\cup\{\varnothing\},\delta_{j})$ complete. While this is true for $j\in\{1,d\}$ (see [Florian-1989, Vitale-1985] for the case $j=1$), as we show below, it is still false for intermediate $j\in\{2,\ldots,d-1\}$.
Theorem 1.5.
Throughout the paper, we use $c,C,C_{1},C_{2}$, etc. to denote positive absolute constants whose value may change from line to line. If a positive constant depends on some parameter, the dependence will always be made explicit, e.g., $c(d)$ is a constant that depends only on $d$.
@@ 2. Proofs
and we integrate this bound over $H$ (the projection is 11-Lipschitz on $E$). Multiplying both sides of this inequality by the fixed factor $\big[\!\begin{smallmatrix}d\\ j\end{smallmatrix}\!\big]$ and setting $C(d,j):=2\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\operatorname{vol}_{j-1}(B_{j-1})$, we obtain (4).
@@
U.S.A.
E-mail address: hoehnersd@longwood.edu

### v2

@@ Abstract.
The purpose of this note is to derive certain basic, but previously unrecorded, topological properties of the intrinsic volume metrics $\delta_{1},\ldots,\delta_{d}$ on the space of convex bodies in $\mathbb{R}^{d}$. Our main results show that for every $2\leq j\leq d-1$, the topology induced by $\delta_{j}$ does not control the Hausdorff metric on the class of $j$-dimensional convex bodies; in particular, the condition $\delta_{j}(K_{n},K)\to 0$ does not imply uniform boundedness in the ambient space. Furthermore, for every $2\leq j\leq d-1$, the metric space $(K_{j}^{d},\delta_{j})$ is incomplete, and remains incomplete even after adjoining the empty set.
Our main results demonstrate that the intrinsic volume metric behaves in a fundamentally different way from the familiar Hausdorff and symmetric difference metrics. We describe the geometric mechanism that produces these phenomena and discuss implications for geometric tomography, metric stability theory and integral geometry.
@@ 1. Introduction and main results
Metrics on the space of convex bodies are indispensable tools throughout convex geometry, geometric analysis and geometric tomography. The Hausdorff metric governs compactness phenomena and the Blaschke selection principle; the Banach–Mazur distance underlies asymptotic classification questions; and the symmetric difference metric is fundamental in shape optimization, stochastic geometry and random approximation theory.
The present paper concerns another natural family of metrics: those induced by the intrinsic volumes. For an integer $d\geq 2$ and a fixed $j\in\{1,\ldots,d\}$, the Kubota formula representation of the $j$th intrinsic volume of a convex body $K\subset\mathbb{R}^{d}$ is
$V_{j}(K)=\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\int_{\mathrm{Gr}(d,j)}\mathrm{vol}_{j}(P_{H}K)\,d\nu_{j}(H).$
@@
Here and throughout the paper, $\operatorname{Gr}(d,j)$ is the Grassmannian manifold of all $j$-dimensional subspaces of $\mathbb{R}^{d}$, $\nu_{j}$ is the unique Haar probability measure on $\operatorname{Gr}(d,j)$, $P_{H}$ denotes the orthogonal projection of $\mathbb{R}^{d}$ into the subspace $H\in\operatorname{Gr}(d,j)$, and
$\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]:=\binom{d}{j}\frac{\operatorname{vol}_{d}(B_{d})}{\operatorname{vol}_{j}(B_{j})\operatorname{vol}_{d-j}(B_{d-j})}$
is the flag coefficient of Klain and Rota [Klain-Rota], where $B_{m}$ is the Euclidean unit ball in $\mathbb{R}^{m}$.
In view of Kubota’s formula, for convex bodies $K,L\subset\mathbb{R}^{d}$ and $j\in\{1,\ldots,d\}$, Besau and Hoehner [BH-2024] recently defined the $j$th intrinsic volume metric by
(1) $\delta_{j}(K,L)=\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\int_{\mathrm{Gr}(d,j)}\operatorname{vol}_{j}\left((P_{H}K)\triangle(P_{H}L)\right)d\nu_{j}(H).$
This metric measures the distance between two convex bodies via the average $j$-dimensional discrepancy of their orthogonal projections. Note that if $j=d$, then (1) reduces to the symmetric difference metric.
Projection-based distances of the form (1) arise naturally in the study of projection inequalities, Urysohn-type inequalities, stability problems, and reconstruction in geometric tomography. In applied contexts, $\delta_{j}$ quantifies differences between shapes when only lower-dimensional data are available, a situation which arises frequently in computerized tomography, stereology, and the analysis of high-dimensional data through random projections. In [BH-2024, BHK, HSW-2025], this metric was used to study the asymptotic best approximation of convex bodies by polytopes.
1.1. Motivation
A longstanding theme in metric geometry is to understand how different metrics interact: when they are equivalent, when they control one another, and when they yield fundamentally different topologies. In general, comparing the various metrics on convex bodies is a fundamental question that has been studied extensively; see, for example, [BH-2024, Florian-1989, Jin-Leng-Guo-2014-orlicz-metric, Groemer2000, ShephardWebster-1965, Vitale-1985].
With the recent introduction of the intrinsic volume metric in [BH-2024], it is thus natural to further investigate how this new metric compares to the other well-studied metrics, such as the Hausdorff and symmetric difference metric. Despite the central role of intrinsic volumes and their projection representations in convex geometry, many of the basic topological properties of the metrics $\delta_{j}$ have not yet been studied. The aim of this paper is to close these gaps while identifying several structural features that distinguish $\delta_{j}$ sharply from the Hausdorff and symmetric difference metrics.
Furthermore, understanding the topology induced by the intrinsic volume metric is important for several reasons, especially as it relates to applications:
  • Stability of geometric inequalities. Many inequalities involving intrinsic volumes, including Urysohn-type and Alexandrov–Fenchel inequalities, require sharp stability estimates. Determining whether $\delta_{j}$-convergence controls geometric degeneracy is helpful for formulating meaningful stability results.
  • Geometric tomography. Reconstruction algorithms often rely on comparing projection data. If $\delta_{j}$-convergence does not prevent “unbounded drift” in the Hausdorff metric, this limits identifiability and affects the choice of regularization.
  • Random approximations and stochastic geometry. Models of random polytopes or random projections approximate a target body in low-dimensional marginals. Whether such approximations converge to genuine convex bodies depends on the completeness and compactness properties of $(K_{j}^{d},\delta_{j})$.
1.2. Main results
Let $\mathcal{K}_{j}^{d}$ denote the set of $j$-dimensional convex bodies contained in some $j$-dimensional affine subspace of $\mathbb{R}^{d}$. Our first theorem shows that the intrinsic-volume metrics do not enforce any form of ambient boundedness.
Theorem 1.1 (Failure of uniform boundedness).
@@ 1.2. Main results
This illustrates that $\delta_{j}$ captures only the average projection geometry of a body, not its global embedding in $\mathbb{R}^{d}$. This distinction is crucial in problems where ambient-space control is essential.
Our next results further establish that $\delta_{j}$ behaves quite differently from the Hausdorff and symmetric difference metrics from a metric-geometric point of view.
Theorem 1.2 (Incompleteness).
It is thus natural to ask if appending the empty set to $\mathcal{K}_{j}^{d}$ will make $(\mathcal{K}_{j}^{d}\cup\{\varnothing\},\delta_{j})$ complete. While this is true for $j\in\{1,d\}$ (see [Florian-1989, Vitale-1985] for the special case $j=1$), as we show below, it is still false for intermediate $j\in\{2,\ldots,d-1\}$.
Theorem 1.3 (Incompleteness with $\varnothing$ adjoined).
These negative results have both analytic and structural implications. They rule out direct use of fixed-point theorems, gradient-flow constructions and variational arguments based on completeness, and they clarify the limitations of projection-based metrics in quantifying geometric proximity.
@@
The proofs of the main theorems are given in Section 3. The geometric idea in the proofs is to construct thin cylinders added far away to produce vanishing projection deviation but large Hausdorff drift, while the projection of cylinders inside “good” subspaces produces persistent mass arbitrarily far away, contradicting possible convergence.
2. Background and notation
where $B_{d}$ is the $d$-dimensional Euclidean unit ball centered at the origin. The symmetric difference metric is defined by $d_{S}(K,L)=\operatorname{vol}_{d}(K\triangle L)$, where $\operatorname{vol}_{d}$ denotes the $d$-dimensional volume functional and $K\triangle L=(K\setminus L)\cup(L\setminus K)$. It is well-known that the metric spaces $(\mathcal{K}^{d},d_{H})$ and $(\mathcal{K}_{d}^{d}\cup\{\varnothing\},d_{S})$ are complete. As explained in [BH-2024], in the latter case the empty set is appended to include limits of sequences of convex bodies $\{K_{i}\}\subset\mathcal{K}^{d}_{d}$ which converge in the Hausdorff metric to a set $K_{0}\in\mathcal{K}^{d}$ with $\dim(K_{0})<d$. More specifically, if $K_{i}\to K_{0}\in\mathcal{K}^{d}$ and $\dim(K_{0})<d$, then for such a sequence we have $d_{S}(K_{i},\varnothing)=\operatorname{vol}_{d}(K_{i})\to 0$, so $K_{i}\to\varnothing$ with respect to $d_{S}$. It is also known that on uniformly bounded subsets of $\mathcal{K}_{d}^{d}$, the metrics $d_{H}$ and $d_{S}$ are equivalent. For more background on the Hausdorff and symmetric difference metrics, see the article [ShephardWebster-1965] by Shephard and Webster. For a general reference on convex geometry, we refer the reader to the book [SchneiderBook] of Schneider.
Remark 2.1.
It was conjectured in [BH-2024] that $d_{H}$ and $\delta_{j}$ induce the same topology on $\mathcal{K}_{j}^{d}$. More specifically, in Open Question 2.2 of [BH-2024] they asked the following
Open Question 2.2 (uniform bound).
For the special case $j=d$, this question was answered affirmatively by Shephard and Webster in [ShephardWebster-1965, Lemma 11], and for $j=1$ see [Vitale-1985, Theorem 3] and [Florian-1989, Theorem 2]. Theorem 1.1 gives a negative answer to this question for $2\leq j\leq d-1$.
@@ 3. Proofs
and we integrate this bound over $H$ (the projection is $1$-Lipschitz on $E$). Multiplying both sides of this inequality by the fixed factor $\big[\!\begin{smallmatrix}d\\ j\end{smallmatrix}\!\big]$ and setting $C(d,j):=2\left[\!\!\begin{array}[]{c}d\\ j\end{array}\!\!\right]\operatorname{vol}_{j-1}(B_{j-1})$, we obtain (4).
@@
201 High Street, Farmville, Virginia 23901
E-mail address:
hoehnersd@longwood.edu